I need help on solving linear equations by elimination by promblem is solv ing this equation : 5(2x-y)=x+20
3x+8=4x-2y please help me solve it!!!!! Its hard

Answer by Maths68(1474)

5(2x-y)=x+20..........(1)
3x+8=4x-2y............(2)
Take (1)
5(2x-y)=x+20
10x-5y=x+20
10x-5y-x=20
9x-5y=20.............(3)
Now (2)
3x+8=4x-2y
3x-4x+2y=-8
-x+2y=-8.............(4)
Multiply (4) by 9 and add to (3)
9x-5y=20
-9x+18y=-72
--------------
13y=-52
y=-52/13
y=-4
Put the value of y in (4)
-x+2y=-8
-x+2(-4)=-8
-x-8=-8
-x=-8+8
-x=0
x=0

x,y=0,-4

Check
=====
Put the values of x and y in (1)
5(2x-y)=x+20..........(1)
5(2(0)-(-4))=0+20
5(0+4)=20
5(4)=20
20=20
Put the values of x and y in (2)
3x+8=4x-2y............(2)
3(0)+8=4(0)-2(-4)
0+8=0+8
8=8